A large tank of fish from a hatchery is being delivered to a lake. The hatchery claims that the mean length of fish in the tank is 15 inches, and the standard deviation is 7 inches. A random sample of 53 fish is taken from the tank. Let x be the mean sample length of these fish. What is the probability that x is within 0.5 inch of the claimed population mean? (Round your answer to four decimal places.)
Answer #1

Answer:

Given Data

Mean = 15

SD = 7

Now , no of sample = 53

Sample SD

=\frac{SD}{\sqrt{n}}

=\frac{7}{\sqrt{53}}

= 0.9615

Let the sample mean be x

X is normally distributed with a mean 15 and SD = 0.9615

Converting to standard normal,

z = \frac{(x-15)}{0.9615}

P ( x lies within 0.5 inch of population mean )

= P(14.5<x<15.5)

=P((14.5-15)/0.9615<z<(15.5-15)/0.9615)

=P(-0.52<z<0.52)

=P(z<0.52)-P(z<-0.52)

=P(z<0.52)-[1-P(z<0.52)]

=2P(z<0.52)-1

=2(0.69847)-1

= 0.39694

***Please like it....

It is important to me...

Thank you for supporting me..